OpenStudy (anonymous):

Given D = [5,0,-1_3,2,4_8,-6,7] and E is additive inverse of D, what is E?

jimthompson5910 (jim_thompson5910):

Hint: If E is additive inverse of D, then D+E = 0 D+E-D = 0-D E = -D E = -1*D

OpenStudy (anonymous):

I needed to find the det first right to figure out if it would have a inverse?

jimthompson5910 (jim_thompson5910):

no you just need to multiply everything in matrix D by -1

jimthompson5910 (jim_thompson5910):

no determinants or inverses needed

OpenStudy (anonymous):

Ok hold on I will get back to u with the answer

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

[-5,0,1_-3,-2,-4_-8,-6,-7] ?